OpenStudy (anonymous):

Using substitution, what is the solution of the system? y = 2x+7 y = x-1 a)(2 2/3,1 2/3) b)(8,-9) c)(-8,-9) d)(1,0)

OpenStudy (mathstudent55):

Since the two equations are solved for y, and in the first equation you see that y equals 2x + 7, insert 2x + 7 in for y in the second equation: 2x + 7 = x - 1 Solve for x: x = -8 Now substitute this value of x into second original equation: y = -8 - 1 y = -9 Solution: x = -8, y = -9

OpenStudy (anonymous):

Thank you!!! @mathstudent55!!

OpenStudy (mathstudent55):

wlcm